OpenStudy (calculusxy):

MEDAL!!!! Algebra 1 help... Let f and g be the functions given by f(x) = x^2 and g(x) = x|x|. a. Is there a value of x, such that f(x) + g(x) = -100 ? If so, find x. If not, explain why no such value exists.

OpenStudy (calculusxy):

@agent0smith

OpenStudy (agent0smith):

So basically\[\large x^2 + x \left| x \right|=-100\]

OpenStudy (calculusxy):

Yes...

OpenStudy (calculusxy):

But how would I solve for x?

OpenStudy (agent0smith):

Well clearly x would have to be negative, that way the x|x| would be negative. But the x^2 will always have same value, since x|x| is really equivalent to -x^2, if x is negative. So your equation, when x is negative, really is \[\large x^2 - x^2 = -100\]

OpenStudy (calculusxy):

So this isn't possible. Right?

OpenStudy (agent0smith):

And if x was positive, the equation is \[\large x^2 + x^2 = -100\]

OpenStudy (calculusxy):

The next one is the same thing, but it's now asking if there's a value that will equal to 50.

OpenStudy (calculusxy):

So how would I determine whether to have x|x| as a negative or a positive?

OpenStudy (agent0smith):

You should be able to figure it out, go through everything i said again

OpenStudy (agent0smith):

I gave two versions of the equation...